# ESND: An Embedding-based Framework for Signed Network Dismantling

### v1

@@ Abstract
Network dismantling aims to maximize the disintegration of a network by removing a specific set of nodes or edges and is applied to various tasks in diverse domains, such as cracking down on crime organizations, delaying the propagation of rumors, and blocking the transmission of viruses. Most of the current network dismantling methods are tailored for unsigned networks, which only consider the connection between nodes without evaluating the nature of the relationships, such as friendship/hostility, enhancing/repressing, and trust/distrust. We here propose an embedding-based algorithm, namely ESND, to solve the signed network dismantling problem. The algorithm generally iterates the following four steps, i.e., giant component detection, network embedding, node clustering, and removal node selection. To illustrate the efficacy and stability of ESND, we conduct extensive experiments on six signed network datasets as well as null models, and compare the performance of our method with baselines. Experimental results consistently show that the proposed ESND is superior to the baselines and displays stable performance with the change in the network structure. Additionally, we examine the impact of sign proportions on network robustness via ESND, observing that networks with a high ratio of negative edges are generally easier to dismantle than networks with high positive edges.
keywords:
@@ 1 Introduction
To address the challenge of signed network dismantling, we propose an algorithm named Embedding-based framework for signed network dismantling (ESND), which integrates node embedding [36] and node clustering to achieve rapid disintegration of a signed network. The ESND consists of three main parts that iteratively remove nodes from the network (see Figure 1): First, we perform a signed network embedding algorithm (SiNE) to obtain node embedding vectors that could capture the local and global structure of a signed network. Second, we employ the K-means algorithm to classify the nodes into different clusters. Lastly, the node with the highest degree in the largest cluster is removed from the network. We compare ESND with the baselines on different empirical signed networks and their null models. The results show that ESND could better dismantle a signed network than the baseline methods.
The subsequent sections of this paper are organized as follows: Section 2 details the specifics of our proposed algorithm. Section 3 offers a clear description of the baseline methods. Section 4 introduces the datasets and presents all the experimental results. We summarize our work and discuss future research directions in Section 5.
2 Method
@@
2.1 Giant Connected Component Detection
@@
2.2 Signed Network Embedding (SiNE)
We choose to use a classic signed network embedding method rooted in deep learning, specifically known as SiNE, to obtain embedding vectors for each node. In the subsequent sections, we provide an in-depth description of the three fundamental components of this method, i.e., the establishment of the objective function, the construction of a deep learning network, and the update of the parameters. The formulation of the objective function in SiNE relies on structural balance theory, positing that individuals are more akin to their “friends” than to their “enemies”. We utilize $\mathcal{T}=\{(v_{i},v_{j},v_{k})\mid e_{ij}=1,e_{ik}=-1,v_{i},v_{j},v_{k}\in V\}$ to denote a collection of triplets, where there is a positive connection between $v_{i}$ and $v_{j}$, and a negative connection between $v_{i}$ and $v_{k}$. Hence, it is necessary to allocate a greater similarity to $v_{i}$ and $v_{j}$ compared to that of $v_{i}$ and $v_{k}$. Mathematically, we express the similarity as $f(\mathbf{x}_{i},\mathbf{x}_{j})\geq f(\mathbf{x}_{i},\mathbf{x}_{k})+\epsilon$, where $f$ denotes the similarity function that requires learning, and $\epsilon$ fine-tunes the dissimilarity between the nodes. The higher value of $\epsilon$ makes $v_{i}$ and $v_{j}$ closer and $v_{i}$ and $v_{k}$ farther away in the embedding space. Since the mentioned function is unable to handle cases where the 2-hop networks of nodes only have positive or negative links, and given that positive connections are more prevalent than negative ones in real-world networks, the study introduces a virtual node $v_{0}$. The virtual node is utilized to establish a negative link between $v_{0}$ and the node connected to its 2-hop neighbors only by positive links. Assuming $\mathcal{T}_{0}=\{(v_{i},v_{j},v_{0})\mid e_{ij}=1,e_{i0}=-1\}$ is one of these triplets, we have $f(\mathbf{x}_{i},\mathbf{x}_{j})\geq f(\mathbf{x}_{i},\mathbf{x}_{0})+\epsilon_{0}$, where $\epsilon_{0}$ plays a similar role as $\epsilon$. Consequently, the objective function for signed network embedding is
$\min_{\mathbf{X},\mathbf{x}_{0},\epsilon}\frac{1}{T}\left[\begin{array}[]{c}\sum\limits_{\left(\mathbf{x}_{i},\mathbf{x}_{j},\mathbf{x}_{k}\right)\in\mathcal{T}}\max\left(0,f\left(\mathbf{x}_{i},\mathbf{x}_{k}\right)+\epsilon-f\left(\mathbf{x}_{i},\mathbf{x}_{j}\right)\right)\\ +\sum\limits_{\left(\mathbf{x}_{i},\mathbf{x}_{j},\mathbf{x}_{0}\right)\in\mathcal{T}_{0}}\max\left(0,f\left(\mathbf{x}_{i},\mathbf{x}_{0}\right)+\epsilon_{0}-f\left(\mathbf{x}_{i},\mathbf{x}_{j}\right)\right)\end{array}\right]$ (1)
@@ 2.2 Signed Network Embedding (SiNE)
The optimization of the objective function is carried out to acquire nonlinear embedding vectors for nodes within signed networks. Within the SiNE framework, the function $f$ and the parameter set $\phi$ in the objective function are defined through the construction of a deep neural network. The framework consists of two layers of deep neural networks, where $\mathbf{W}^{11}$ and $\mathbf{W}^{12}$ are the weights of the first hidden layer, and $\mathbf{b}^{1}$ is the bias. The specific output form of the first layer is as follows:
@@
Similarly, the outputs of the first layer, $\mathbf{z}^{11}$ and $\mathbf{z}^{12}$, serve as inputs of the second layer. The specific structure of the output of the second hidden layer is expressed as $\mathbf{z}^{21}=\tanh(\mathbf{W}^{2}\mathbf{z}^{11}+\mathbf{b}^{2})$ and $\mathbf{z}^{22}=\tanh(\mathbf{W}^{2}\mathbf{z}^{12}+\mathbf{b}^{2})$, where $\mathbf{W}^{2}$ represents the weight of the second-layer network, and $\mathbf{b}^{2}$ denotes the bias. Thus, the final output of the deep neural network determines the nonlinear function $f$ used to evaluate node similarity in the objective function, which can be expressed as
@@ 4 Experiments
4.1 Datasets
We select six real-world datasets that can be constructed as signed networks to evaluate the performance of our method. Specifically, Bitcoinalpha and Bitcoinotc are data sourced from SNAP  [^1] https://snap.stanford.edu/data/, illustrating the trust networks between users participating in Bitcoin transactions. Due to transactional anonymity in Bitcoin transactions, users provide positive and negative ratings to signify trust (positive) or distrust (negative) relationships. WikiVote represents the voting network to select Wikipedia administrators [^2] https://doi.org/10.6084/m9.figshare.12152628. The eligibility of the users for administration is determined through voting, with the edges denoting voting interactions, i.e., positive signs indicate support while negative signs indicate opposition. Slashdot is a notable technology news site where users comment and share technology-related information [^3] https://www.aminer.cn/data-sna. Positive and negative signs in the dataset denote friendly or adversarial relationships between users. Reddit captures connections between users in diverse sub-communities, reflecting positive or negative sentiment in shared content across online accounts2. Epinions constitutes a trust network among users on a product review website, with positive and negative signs indicating trust or distrust relationships between user connections3. We show the topological information of these signed networks in Table 1, including the number of nodes ($N$), the number of edges ($M$), the number of positive edges ($E^{+}$), the number of negative edges ($E^{-}$), the average degree of nodes ($\left\langle k\right\rangle$) and the clustering coefficient ($C$). The table shows that all the signed networks have more positive edges than negative ones.
Table 1: Topological information of the signed networks, in which $N$ denotes the number of nodes; $M$ represents the number of edges; $\left|E^{+}\right|$ and $\left|E^{-}\right|$ indicate the number of positive and negative edges, respectively. The values in parentheses represent the proportions of positive and negative edges in the network; $\left\langle{k}\right\rangle$ denotes the average degree, and $C$ signifies the average clustering coefficient.
@@ 4.1 Datasets
4.2 Performance Evaluation Metric
@@ 4.3 Parameter Analysis
4.4 Experimental Results
4.4.1 Results on Real Signed Networks
We compare the performance of the ESND with the selected baselines on the six signed networks, where the results are given in Figure 3 and Table 2. In Figure 3, the horizontal axis ($q$) represents the proportion of nodes removed, while the vertical axis ($S(qN)$) corresponds to the fraction of nodes in the GCC after removing $q$ fraction of nodes. For a fixed value of $q$, the smaller value of $S(qN)$ indicates that the dismantling method is more effective in dismantling the corresponding signed network than other methods. The values in Table 2 reveal the area under each curve (AUC) in Figure 3, with a smaller value indicating the better performance of the corresponding dismantling method. The experimental results show that the robustness of these real signed networks is notably different, with some of them demonstrating fast network collapse with only a small fraction of nodes being removed, such as Slashdot and Epinions, while the remaining ones are more robust. For example, WikiVote and Reddit networks necessitate approximately 40% removal to attain complete decomposition for most of the dismantling methods. In addition, ESND outperforms all baseline methods in dismantling most signed networks, particularly when we remove a large fraction of nodes. In dismantling an unsigned network, normally the betweenness can outperform the other methods in most cases[45]. However, it performs second best in most cases in dismantling signed networks, which reveals that considering the topology deduced by the signs is important in dismantling a signed network. Moreover, various centrality methods exhibit varying performances across different datasets, including Closeness, K-shell, PN, TE, SPR, and SE. The efficacy of these methods is closely related to the specific structures of the networks. In contrast, ESND consistently achieves optimal network dismantling results across diverse datasets, showing its stability and effectiveness.
@@ 4.4.1 Results on Real Signed Networks
Table 2: Area under each curve (AUC) of each curve in Figure 3. The best performance is highlighted in bold.
@@
4.4.2 Results on Signed Null Models
To delve deeper into the potential influence of factors such as network topology and signs on ESND and their consequent impact on variations in network dismantling outcomes, three distinct null models for signed networks[46] were constructed in six datasets. These null models exhibit certain properties akin to the original networks. Detailed introductions to them are provided below.
  • 1.
    Rewire: Given a signed network, the model exchanges edges between nodes while keeping the degree of each node unchanged. In this null model, both the topological structure of the network and the positive and negative degrees of each node undergo alterations.
  • 2.
    Signed rewire: Initially, this model constructs positive and negative edge subgraphs from the original network. Subsequently, it reconnects the edges within these two subgraphs, preserving the positive and negative degree of each node. The process ends by merging the two rewired subgraphs to establish the null model. In this model, the positive and negative degrees of nodes remain unchanged, while the network structure undergoes modifications.
  • 3.
    Sign shuffle: In this model, the topological structure of the network is preserved by randomly selecting one positive edge and one negative edge and exchanging their signs, thereby introducing random perturbations to the signs of the edges.
@@ 4.4.2 Results on Signed Null Models
We conduct network dismantling on these null models generated by each of the six real-world signed networks, the specific experimental results are illustrated in Figure 5, where each row corresponds to a signed network and its null models. The horizontal axis denotes various network dismantling methods, while the vertical axis represents the evaluation metrics $R$ for these methods in various signed networks and null models. The results show that each of the dismantling methods demonstrates generally consistent performance in network dismantling across both the original network and three null models within the same dataset. This suggests that modifications to the topological properties and sign distribution of the signed network do not significantly affect the efficacy of these methods. More importantly, ESND consistently attains superior dismantling performance compared to these baselines across these varied null models, emphasizing the stability of ESND as an effective method for dismantling networks.
5 Conclusion
@@
5.0.1 Impact of the signs on network robustness
In Figure 6, the x-axis indicates the ratio of negative edges ($\frac{|E^{-}|}{M}$) in each of the networks, and the y-axis shows the $R$ values, revealing the robustness of the corresponding networks. Although the WS network has a higher value of $R$ (indicating more robustness) for a low value of $\frac{|E^{-}|}{M}$ compared to ER and BA, it is easier to dismantle when $\frac{|E^{-}|}{M}>0.4$. Meanwhile, we observe that as the ratio of negative edges increases for a relatively small value of $\frac{|E^{-}|}{M}$ ($\frac{|E^{-}|}{M}<0.4$ for WS, $\frac{|E^{-}|}{M}<0.7$ for ER and BA), the robustness of the networks is relatively stable. However, for a large value of $\frac{|E^{-}|}{M}$, the networks can easily be dismantled, with the value of $R$ decreasing with increasing $\frac{|E^{-}|}{M}$. This suggests that increasing the proportion of positive edges in the network contributes to enhancing its robustness. This observation aligns with real-world scenarios. In a social network where negative edges dominate, signifying antagonistic relationships between individuals, the network is naturally more vulnerable. In general, ER and BA networks are more resilient than the WS network with increasing $\frac{|E^{-}|}{M}$.
Certain aspects of this study warrant further attention. Future work could explore the following aspects: a) in this work, we treat signed networks as undirected networks. However, most of the signed networks in the real world contain directionality. Therefore, how to efficiently dismantle a directed signed network remains a topic worth investigating. b) We confine ourselves to using unsigned network connectivity, i.e., the fraction of nodes in the giant component, to evaluate the performance of a dismantling method. Future work could also propose performance evaluation methods regarding the sign nature of a network. c) Most of the existing research on network dismantling focuses on removing critical nodes. Devising methods that could identify critical edges in signed networks to achieve rapid network decomposition is also an
6 Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
7 Acknowledgement

### v2

@@ Abstract
Network dismantling aims to maximize the disintegration of a network by removing a specific set of nodes or edges and is applied to various tasks in various domains, such as cracking down on crime organizations, delaying the propagation of rumors, and blocking the transmission of viruses. Most of the current network dismantling methods are tailored for unsigned networks, which only consider the connection between nodes without evaluating the nature of the relationships, such as friendship/hostility, enhancing/repressing, and trust/distrust. We here propose an embedding-based algorithm, namely ESND, to solve the signed network dismantling problem. The algorithm generally iterates the following four steps, i.e., giant component detection, network embedding, node clustering, and removal node selection. To illustrate the efficacy and stability of ESND, we conduct extensive experiments on six signed network datasets as well as null models, and compare the performance of our method with baselines. Experimental results consistently show that the proposed ESND is superior to the baselines and displays stable performance with the change in the network structure. Additionally, we examine the impact of sign proportions on network robustness via ESND, observing that networks with a high ratio of negative edges are generally easier to dismantle than networks with high positive edges.
Index Terms:
@@ I Introduction
To address the challenge of signed network dismantling, we propose an algorithm named the Embedding-based framework for signed network dismantling (ESND), which integrates node embedding [36] and node clustering to achieve rapid disintegration of a signed network. The ESND consists of three main parts that iteratively remove nodes from the network (see Figure 1): First, we perform a signed network embedding algorithm (SiNE) to obtain node embedding vectors that could capture the local and global structure of a signed network. Second, we employ the K-means algorithm to classify the nodes into different clusters. Lastly, the node with the highest degree in the largest cluster is removed from the network. We compare ESND with the baselines on different empirical signed networks and their null models. The results show that ESND could better dismantle a signed network than the baseline methods.
The subsequent sections of this paper are organized as follows. Section II details the specifics of our proposed algorithm. Section III offers a clear description of the baseline methods. Section IV introduces the datasets and presents all the experimental results. We summarize our work and discuss future research directions in Section V.
II Methods
@@
II-A Giant Connected Component Detection
@@
II-B Signed Network Embedding (SiNE)
We choose to use a classic signed network embedding method rooted in deep learning, specifically known as SiNE, to obtain embedding vectors for each node. In the subsequent sections, we provide an in-depth description of the three fundamental components of this method, i.e., the establishment of the objective function, the construction of a deep learning network, and the update of the parameters. The formulation of the objective function in SiNE is based on structural balance theory, positing that individuals are more like their “friends” than their “enemies”. We utilize $\mathcal{T}=\{(v_{i},v_{j},v_{k})\mid e_{ij}=1,e_{ik}=-1,v_{i},v_{j},v_{k}\in V\}$ to denote a collection of triplets, where there is a positive connection between $v_{i}$ and $v_{j}$, and a negative connection between $v_{i}$ and $v_{k}$. Hence, it is necessary to allocate a greater similarity to $v_{i}$ and $v_{j}$ compared to $v_{i}$ and $v_{k}$. Mathematically, we express the similarity as $f(\mathbf{x}_{i},\mathbf{x}_{j})\geq f(\mathbf{x}_{i},\mathbf{x}_{k})+\epsilon$, where $f$ denotes the similarity function that requires learning, and $\epsilon$ fine-tunes the dissimilarity between the nodes. The higher value of $\epsilon$ makes $v_{i}$ and $v_{j}$ closer and $v_{i}$ and $v_{k}$ farther away in the embedding space. Since the mentioned function is unable to handle cases where 2-hop networks of nodes only have positive or negative links, and given that positive connections are more prevalent than negative ones in real-world networks, the study introduces a virtual node $v_{0}$. The virtual node is utilized to establish a negative link between $v_{0}$ and the node connected to its 2-hop neighbors only by positive links. Assuming $\mathcal{T}_{0}=\{(v_{i},v_{j},v_{0})\mid e_{ij}=1,e_{i0}=-1\}$ is one of these triplets, we have $f(\mathbf{x}_{i},\mathbf{x}_{j})\geq f(\mathbf{x}_{i},\mathbf{x}_{0})+\epsilon_{0}$, where $\epsilon_{0}$ plays a similar role as $\epsilon$. Consequently, the objective function for signed network embedding is
$\min_{\mathbf{X},\mathbf{x}_{0},\epsilon}$ $\frac{1}{T}\left[\sum\limits_{\left(\mathbf{x}_{i},\mathbf{x}_{j},\mathbf{x}_{k}\right)\in\mathcal{T}}\max\left(0,f\left(\mathbf{x}_{i},\mathbf{x}_{k}\right)+\epsilon-f\left(\mathbf{x}_{i},\mathbf{x}_{j}\right)\right)\right.$ (1)
$\left.+\sum\limits_{\left(\mathbf{x}_{i},\mathbf{x}_{j},\mathbf{x}_{0}\right)\in\mathcal{T}_{0}}\max\left(0,f\left(\mathbf{x}_{i},\mathbf{x}_{0}\right)+\epsilon_{0}-f\left(\mathbf{x}_{i},\mathbf{x}_{j}\right)\right)\right]$
@@ II-B Signed Network Embedding (SiNE)
The optimization of the objective function is carried out to acquire nonlinear embedding vectors for nodes within signed networks. Within the SiNE framework, the function $f$ and the parameter set $\phi$ in the objective function are defined through the construction of a neural network. The framework consists of two layers of neural networks, where $\mathbf{W}^{11}$ and $\mathbf{W}^{12}$ are the weights of the first hidden layer, and $\mathbf{b}^{1}$ is the bias. The specific output form of the first layer is as follows:
@@
Similarly, the outputs of the first layer, $\mathbf{z}^{11}$ and $\mathbf{z}^{12}$, serve as inputs of the second layer. The specific structure of the output of the second hidden layer is expressed as $\mathbf{z}^{21}=\tanh(\mathbf{W}^{2}\mathbf{z}^{11}+\mathbf{b}^{2})$ and $\mathbf{z}^{22}=\tanh(\mathbf{W}^{2}\mathbf{z}^{12}+\mathbf{b}^{2})$, where $\mathbf{W}^{2}$ represents the weight of the second-layer network, and $\mathbf{b}^{2}$ denotes the bias. Thus, the final output of the neural network determines the nonlinear function $f$ used to evaluate node similarity in the objective function, which can be expressed as
@@ IV Experiments
IV-A Datasets
We select six real-world datasets that can be constructed as signed networks to evaluate the performance of our method. Specifically, Bitcoinalpha and Bitcoinotc are data sourced from SNAP  [^1] https://snap.stanford.edu/data/, illustrating the trust networks between users participating in Bitcoin transactions. Due to transactional anonymity in Bitcoin transactions, users provide positive and negative ratings to signify trust (positive) or distrust (negative) relationships. WikiVote represents the voting network to select Wikipedia administrators [^2] https://doi.org/10.6084/m9.figshare.12152628. The eligibility of the users for administration is determined through voting, with the edges denoting voting interactions, i.e., positive signs indicate support while negative signs indicate opposition. Slashdot is a notable technology news site where users comment and share technology-related information [^3] https://www.aminer.cn/data-sna. Positive and negative signs in the dataset denote friendly or adversarial relationships between users. Reddit captures connections between users in diverse sub-communities, reflecting positive or negative sentiment in shared content across online accounts2. Epinions constitutes a trust network among users on a product review website, with positive and negative signs indicating trust or distrust relationships between user connections3. We show the topological information of these signed networks in Table I, including the number of nodes ($N$), the number of edges ($M$), the number of positive edges ($E^{+}$), the number of negative edges ($E^{-}$), the average degree of nodes ($\left\langle k\right\rangle$) and the clustering coefficient ($C$). The table shows that all the signed networks have more positive edges than negative ones.
TABLE I: Topological information of the signed networks, in which $N$ denotes the number of nodes; $M$ represents the number of edges; $\left|E^{+}\right|$ and $\left|E^{-}\right|$ indicate the number of positive and negative edges, respectively. The values in parentheses represent the proportions of positive and negative edges in the network; $\left\langle{k}\right\rangle$ denotes the average degree, and $C$ signifies the average clustering coefficient.
@@ IV-A Datasets
IV-B Performance Evaluation Metric
@@ IV-C Parameter Analysis
IV-D Experimental Results
IV-D1 Results on Real Signed Networks
We compare the performance of the ESND with the selected baselines on the six signed networks, where the results are given in Figure 3 and Table II. In Figure 3, the horizontal axis ($q$) represents the proportion of nodes removed, while the vertical axis ($S(qN)$) corresponds to the fraction of nodes in the GCC after removing $q$ fraction of nodes. For a fixed value of $q$, the smaller value of $S(qN)$ indicates that the dismantling method is more effective in dismantling the corresponding signed network than other methods. The values in Table II reveal the area under each curve (AUC) in Figure 3, with a smaller value indicating the better performance of the corresponding dismantling method. The experimental results show that the robustness of these real signed networks is notably different, with some of them demonstrating fast network collapse with only a small fraction of nodes being removed, such as Slashdot and Epinions, while the remaining ones are more robust. For example, WikiVote and Reddit networks necessitate approximately 40% removal to attain complete decomposition for most of the dismantling methods. In addition, ESND outperforms all baseline methods in dismantling most signed networks, particularly when we remove a large fraction of nodes. In dismantling an unsigned network, normally the betweenness can outperform the other methods in most cases[45]. However, it performs second best in most cases in dismantling signed networks, which reveals that considering the topology deduced by the signs is important in dismantling a signed network. Moreover, various centrality methods exhibit varying performances across different datasets, including Closeness, K-shell, PN, TE, SPR, and SE. The efficacy of these methods is closely related to the specific structures of the networks. In contrast, ESND consistently achieves optimal network dismantling results across diverse datasets, showing its stability and effectiveness.
@@ IV-D1 Results on Real Signed Networks
TABLE II: Area under each curve (AUC) of each curve in Figure 3. The best performance is highlighted in bold.
@@
IV-D2 Results on Null Models of the Signed Network
To delve deeper into the potential influence of factors such as network topology and signs on ESND and their consequent impact on variations in network dismantling outcomes, three distinct null models for signed networks[46] were constructed in six datasets. We illustrate examples of the null models in Figure 5, in which they preserve certain properties of the original network. Detailed descriptions of them are given below.
Figure 5: Toy examples of null models of a signed network. Solid lines represent positive edges, and dotted lines represent negative edges.
  • Sign shuffle: In this model, the topological structure of the network is preserved by randomly selecting one positive edge and one negative edge and exchanging their signs, but the positive and negative degrees of each node will change. Taking node $v_{1}$ in Figure 5a and b as an example, the degree of node $v_{1}$ is preserved, but the positive degree and negative degree of node $v_{1}$ change from $\{2,0\}$ to $\{1,1\}$ via the sign shuffle model.
  • Signed rewire: Initially, two subgraphs containing only positive or negative edges are constructed from the original network. Subsequently, the edges are rewired within each subgraph, which could preserve the positive and negative degrees of the nodes. The process ends by merging the two rewired subgraphs to establish the null model. In this model, the positive and negative degrees of the nodes remain the same as in the original network, while the network structure is changed. Figure 5c demonstrates the generation of a signed rewire null model. For example, we disconnect the edges $(v_{1},v_{2})$, $(v_{5},v_{6})$ and form new edges $(v_{1},v_{5})$, $(v_{2},v_{6})$ but keep the positive and negative degree of each node.
  • Rewire: The model exchanges edges between nodes while keeping the degree of each node unchanged. In this null model, both the topological structure of the network and the positive and negative degrees of each node undergo alterations. In Figure 5d, we show that the degree, positive degree, and negative degree of each node are changed through the random rewiring process of the rewire model.
@@ IV-D2 Results on Null Models of the Signed Network
We perform network dismantling on these null models generated by each of the six signed real-world networks, the specific experimental results are illustrated in Figure 6. The horizontal axis denotes the original signed network and their corresponding null models, while the vertical axis represents the evaluation metrics $R$ for various network dismantling methods applied to these signed networks and null models. The results show that each of the dismantling methods demonstrates generally consistent performance in network dismantling across both the original network and three null models within the same dataset. This suggests that modifications to the topological properties and sign distribution of the signed network do not significantly affect the efficacy of these methods. More importantly, ESND consistently attains superior dismantling performance compared to these baselines across these varied null models (as shown by the red diamonds in the figures), emphasizing the stability of ESND as an effective method for dismantling networks.
IV-D3 Impact of the signs on network robustness
In Figure 7, the x-axis indicates the ratio of negative edges ($\frac{|E^{-}|}{M}$) in each of the networks, and the y-axis shows the $R$ values, revealing the robustness of the corresponding networks. Although the WS network has a higher value of $R$ (indicating more robustness) for a low value of $\frac{|E^{-}|}{M}$ compared to ER and BA, it is easier to disassemble when $\frac{|E^{-}|}{M}>0.4$. Meanwhile, we observe that as the ratio of negative edges increases for a relatively small value of $\frac{|E^{-}|}{M}$ ($\frac{|E^{-}|}{M}<0.4$ for WS, $\frac{|E^{-}|}{M}<0.7$ for ER and BA), the robustness of the networks is relatively stable. However, for a large value of $\frac{|E^{-}|}{M}$, the networks can easily be dismantled, with the value of $R$ decreasing with increasing $\frac{|E^{-}|}{M}$. This suggests that increasing the proportion of positive edges in the network contributes to enhancing its robustness. This observation aligns with real-world scenarios. In a social network where negative edges dominate, signifying antagonistic relationships between individuals, the network is naturally more vulnerable. In general, ER and BA networks are more resilient than the WS network with increasing $\frac{|E^{-}|}{M}$.
V Conclusion
@@
Certain aspects of this study warrant further attention. Future work could explore the following aspects: a) in this work, we treat signed networks as undirected networks. However, most of the signed networks in the real world contain directionality. Therefore, how to efficiently dismantle a directed signed network remains a topic worth investigating. b) We confine ourselves to using unsigned network connectivity, i.e., the fraction of nodes in the giant component, to evaluate the performance of a dismantling method. Future work could also propose performance evaluation methods regarding the sign nature of a network. c) Most of the existing research on network dismantling focuses on removing critical nodes. Devising methods that could identify critical edges in signed networks to achieve rapid network decomposition is also an interesting direction that deserves further exploration.
Acknowledgment